# Graph Condensation: A Survey

### v1

@@ Abstract
The burgeoning volume of graph data poses significant challenges in storage, transmission, and particularly the training of graph neural networks (GNNs). To address these challenges, graph condensation (GC) has emerged as an innovative solution. GC focuses on synthesizing a compact yet highly representative graph, on which GNNs can achieve performance comparable to trained on the large original graph. The notable efficacy of GC and its broad prospects have garnered significant attention and spurred extensive research. This survey paper provides an up-to-date and systematic overview of GC, organizing existing research into four categories aligned with critical GC evaluation criteria: effectiveness, generalization, fairness, and efficiency. To facilitate an in-depth and comprehensive understanding of GC, we examine various methods under each category and thoroughly discuss two essential components within GC: optimization strategies and condensed graph generation. Additionally, we introduce the applications of GC in a variety of fields, and highlight the present challenges and novel insights in GC, promoting advancements in future research.
@@ 1 Introduction
Graph data is extensively utilized across a diverse range of domains, owing to its capability to represent complex structural relationships among various entities in the real world Wu et al. (2022); Bo et al. (2023). Notable applications include, but are not limited to, social networks Li et al. (2018); Jung et al. (2012), chemical molecular structures Guo et al. (2022), transportation systems Rahmani et al. (2023), and recommender systems Zheng et al. (2023a); Yu et al. (2023a). However, the exponential growth of data volume in these applications makes the management and processing of large-scale graph data a complex and challenging task. A particularly demanding issue is the considerable computational load associated with training graph neural networks (GNNs) on large-scale graph datasets Hamilton et al. (2017). These challenges are more pronounced in scenarios that require training multiple GNNs, such as hyper-parameter optimization Ding et al. (2022) and neural architecture search Oloulade et al. (2021). This underscores the growing importance for developing more efficient and effective methodologies for processing large-scale graph data.
Figure 1: Taxonomy and evaluation criteria of graph condensation.
@@
A direct solution to these challenges is graph reduction, where appropriate simplification of graphs can not only expedite graph algorithms but also improve the efficiency of storage, transformation and retrieval processes for diverse graph analysis tasks. Early explorations into graph reduction primarily focused on graph sparsification Spielman and Srivastava (2008) and coarsening Loukas and Vandergheynst (2018) techniques. These methods aim to reduce graph size by eliminating redundant edges and merging similar nodes. While effective in maintaining the essential characteristics of large graphs, these approaches have notable limitations. They predominantly rely on heuristic methods, such as the largest principal eigenvalues Loukas and Vandergheynst (2018) or pairwise distances Bravo Hermsdorff and Gunderson (2019), which can result in poor generalization across different GNNs architecture or downstream tasks. Consequently, the focus of research has shifted towards more sophisticated and adaptable methods for graph reduction. A leading development is graph condensation (GC) (a.k.a. graph distillation). GC Jin et al. (2022b) aims to synthesizes a small but highly informative graph dataset, encompassing both node attribute and topology structure, to effectively represent large original graph. Through task-driven optimization, models trained on these small condensed datasets can achieve performance comparable to those trained on the large original graphs. For instance, GCond Jin et al. (2022b) can condense the graph of Flickr dataset to just 0.1% of its original size, while some certain GNNs models trained on this small graph manage to retain 99.8% of the original test accuracy. Given the notable efficacy of GC, it has been soon applied to a variety of applications Gao et al. (2024); Liu et al. (2023b); Ding et al. (2022) and has sparked a proliferation of follow-up research Zheng et al. (2023c); Xu et al. (2023), further broadening the reach and deepening the impact of GC.
@@
Although significant strides have been made to enhance the effectiveness of GC methods, as a data-centric approach, the evaluation of GC extends beyond mere effectiveness, motivating a broader exploration of additional criteria in assessing the performance of GC. One critical aspect is the condensed graph’s ability to generalize across multiple GNN architectures and various downstream tasks Jin et al. (2022b). A GC method with robust generalization is particularly beneficial in scenarios that demand training heterogeneous models or handling diverse tasks Yang et al. (2023a). In addition, it is essential for the condensed graph to maintain unbiased representations compared to the original graph, which is particularly crucial in fields sensitive to bias, such as human resources Bourmpoulias et al. (2023), criminal justice Lettieri et al. (2023), and financial services Wang et al. (2021). Beyond these considerations, the scalability of GC also emerges as a key factor in practical applications Gao et al. (2023b). An efficient GC method enables rapid adaption in life-long graph learning scenarios, effectively handling the continuous growth and dynamic changes of graph data. These criteria play a pivotal role in the evaluation of GC and simultaneously serve as guiding principles for research motivations. Thus, we categorize methods based on these criteria as shown in Figure 1, to provide a structured taxonomy for a comprehensive field survey. This taxonomy can help researchers in comprehensively understanding the diverse motivations driving GC methods, thereby facilitating a clearer path toward achieving their specific objectives.
@@
While recent surveys on related topics including dataset distillation Geng et al. (2023); Yu et al. (2023b); Sachdeva and McAuley (2023); Lei and Tao (2024), data-centric graph learning Yang et al. (2023b); Zheng et al. (2023b); Shabani et al. (2024), as well as graph neural network acceleration Zhang et al. (2023) have touched upon GC, they summarize GC as a peripheral application or future direction, lacking systematic updates on the latest research and a holistic view. Moreover, existing dataset distillation surveys focus on detailed optimization strategies, missing high-level perspectives and design objectives analysis. This hinders researchers from understanding novel GC for varying contexts. Recognizing the rapid evolution and growing significance of GC, there is an urgent need for a systematic survey that encapsulates the latest advancements and motivations behind methodological designs. In summary, our contributions are as follows:
@@
  • We introduces a systematic taxonomy for GC, categorizing existing methods into four distinct groups based on key criteria: effectiveness, generalization, fairness, and efficiency. This taxonomy not only structures existing knowledge, but captures the underlying motivations driving GC research.
@@
  • We provide a comprehensive and up-to-date review of the latest advancements in GC. Additionally, we delve into two fundamental aspects: the optimization strategies and condensed graph generation, facilitating an in-depth understanding of GC techniques.
@@
  • We discuss a variety of practical GC applications and shed light on current challenges and future directions in this field. This discussion aims to guide researchers who are seeking to advance GC further.
@@
The structure of this survey is outlined as follows: Section 2 delves into the definition, design criteria, and taxonomy of GC. Following this, Section 3 discusses the optimization strategies employed in GC methods. Section 4 provides a thorough investigation of each category of GC, as aligned with our taxonomy. In Section 5, we extensively discuss the methods involved in condensed graph generation. Section 6 is dedicated to summarizing the various applications of GC. Finally, we present the challenges and future directions in this field in Section 7.
@@ 2 Definition and Taxonomy
2.1 Problem Definition
@@
Given a graph dataset $\mathcal{T}=\left(\mathcal{V},\mathcal{E}\right)$ with $|\mathcal{V}|=N$ nodes and $|\mathcal{E}|=M$ edges, its adjacency matrix, node feature matrix and task specific label are denoted as ${\mathbf{A}}$, $\mathbf{X}$ and $\mathbf{Y}$, respectively [^2] We unify definitions for node-level and graph-level graph datasets.. Graph condensation aims to create a small condensed graph dataset $\mathcal{S}=\left(\mathcal{V}^{\prime},\mathcal{E}^{\prime}\right)$ with $|\mathcal{V}^{\prime}|=N^{\prime}$ and $N^{\prime}\ll N$. Its adjacency matrix, node feature matrix and task specific label are denoted as ${\mathbf{A}^{\prime}}$, $\mathbf{X}^{\prime}$ and $\mathbf{Y}^{\prime}$, respectively. To facilitate the connection between the original graph $\mathcal{T}$ and the condensed graph $\mathcal{S}$, a relay graph model $f_{\theta}\left(\cdot\right)$, parameterized by ${\theta}$, is employed in the optimization process for encoding both graphs as shown in Figure 2. Then, the graph condensation is formulated as:
$\mathcal{S}=\arg\min_{\mathcal{S}}\mathcal{L}_{cond}\left(f_{\theta}\left({\mathcal{T}},{\mathcal{S}}\right)\right),$ (1)
@@
Figure 2: Graph condensation procedure. The original graph and condensed graph are encoded by relay model $f_{\theta}$ and the condensed graph is optimized according to $\mathcal{L}_{cond}$.
@@ 2.2 Criteria for Graph Condensation
While the accuracy of a specific GNN model, trained on the condensed graph and evaluated with the original graph, is considered the most important metric for evaluating a GC method, there are other criteria that merit consideration. Specifically, we summarize the key criteria of GC as follows:
  • Effectiveness assesses the accuracy of models trained on condensed graphs at various compression rates, using a specified GNN architecture Jin et al. (2022b). It measures the extent to which the condensed graph preserves the essential characteristics of the original graph.
  • Generalization measures the overall accuracy of different GNN architectures trained on the condensed graph across various downstream tasks Yang et al. (2023a). It reflects the condensed graph’s adaptability and utility in diverse contexts.
  • Fairness evaluates the disparity in fairness between models trained on condensed graphs and those trained on original graphs Feng et al. (2023), aiming to prevent bias amplification in the condensed graph and ensure equitable model performance.
  • Efficiency assesses the time required to learn a condensed graph Jin et al. (2022a), focusing on the speed of condensed graph generation and the method’s feasibility for time-sensitive scenarios.
Given these criterion, an ideal GC method should generate a high-quality condensed graph dataset, enabling models with different GNN architectures to achieve results comparable and unbiased to those obtained on the original graph on a wide range of downstream tasks.

### v2

@@ Abstract
The rapid growth of graph data poses significant challenges in storage, transmission, and particularly the training of graph neural networks (GNNs). To address these challenges, graph condensation (GC) has emerged as an innovative solution. GC focuses on synthesizing a compact yet highly representative graph, enabling GNNs trained on it to achieve performance comparable to those trained on the original large graph. The notable efficacy of GC and its broad prospects have garnered significant attention and spurred extensive research. This survey paper provides an up-to-date and systematic overview of GC, organizing existing research into five categories aligned with critical GC evaluation criteria: effectiveness, generalization, efficiency, fairness, and robustness. To facilitate an in-depth and comprehensive understanding of GC, this paper examines various methods under each category and thoroughly discusses two essential components within GC: optimization strategies and condensed graph generation. We also empirically compare and analyze representative GC methods with diverse optimization strategies based on the five proposed GC evaluation criteria. Finally, we explore the applications of GC in various fields, outline the related open-source libraries, and highlight the present challenges and novel insights, with the aim of promoting advancements in future research. The related resources can be found at https://github.com/XYGaoG/Graph-Condensation-Papers.
Index Terms:
Graph Condensation, Data-Centric AI, Graph Representation Learning, Graph Neural Network
@@ 1 Introduction
Graph data is extensively utilized across a diverse range of domains, owing to its capability to represent complex structural relationships among various entities in the real world [1, 2]. Notable applications include, but are not limited to, social networks [3, 4], chemical molecular structures [5], transportation systems [6], and recommender systems [7, 8]. However, the exponential growth of data volume in these applications makes the management and processing of large-scale graph data a complex and challenging task. A particularly demanding aspect is the computational requirements associated with training graph neural networks (GNNs) on large-scale graph datasets [9]. These challenges are more pronounced in scenarios that require training multiple GNNs, such as hyper-parameter optimization [10], continual learning [11] and neural architecture search [12], etc. This underscores the growing importance of developing efficient and effective methodologies for processing large-scale graph data.
Figure 1: An overview for graph condensation. Graph condensation aims to generate small informative graphs such that the models trained on these graphs have similar downstream task performance to those trained on the original graphs. GC methods can be categorised into five classes aligned with critical GC evaluation criteria.
@@
Early attempts to accelerate GNN training predominantly focused on model-centric methods, which aim to design advanced model structures to alleviate the heavy computational burden associated with the message-passing paradigm [13] in GNNs. Specifically, sampling-based methods [9, 14, 15] attempt to minimize the number of nodes involved in message-passing by selectively sampling node neighbors. Linear aggregation methods [16, 17, 18] propagate messages as a pre-processing step to reduce the online training computation. Despite achieving more elaborate model architectures, these model-centric methods suffer from poor generalization and fail to ensure optimal performance across diverse graph datasets and downstream tasks [19]. In light of these limitations, research attention has shifted towards data-centric methods [20], and graph reduction techniques have been proposed to accelerate model training by reducing graph size. Appropriate simplification of graphs can not only expedite graph algorithms but also enhance the efficiency of storage, transformation, and retrieval processes for diverse graph analysis tasks. Early explorations into graph reduction primarily focused on graph sparsification [21] and coarsening [22] techniques. These methods aim to reduce graph size by eliminating redundant edges and merging similar nodes. While effective in maintaining the essential characteristics of large graphs, these approaches have notable limitations. They predominantly rely on heuristic methods, such as the largest principal eigenvalues [22] or pairwise distances [23], which still result in poor generalization across different downstream tasks [24]. Consequently, the emphasis of current research is on developing more generalized and adaptable graph reduction techniques, and the leading development is graph condensation (GC) (a.k.a. graph distillation). As shown in Figure 1, GC [25] aims to synthesize a small but highly informative graph dataset, encompassing both node feature and topology structure, to effectively represent the large original graph. Through task-driven optimization, models trained on these small condensed datasets can achieve performance comparable to those trained on the large original graphs. For instance, GCond [25] can condense the Flickr graph dataset to just 0.1% of its original size, while some certain GNNs models trained on this small graph manage to retain 99.8% of the original test accuracy. Given the notable efficacy of GC, it has soon been applied to a variety of applications [26, 27, 10] and has sparked a proliferation of follow-up research [28, 29], further broadening the reach and deepening the impact of GC.
@@
Although significant strides have been made to enhance the effectiveness of GC methods, their evaluation extends beyond mere effectiveness due to the data-centric focus of these methods. This necessitates a broader exploration of additional criteria for assessing the performance of GC. One critical aspect is the condensed graph’s ability to generalize across multiple GNN architectures and various downstream tasks [25]. A GC method with superior generalization is particularly beneficial in scenarios that demand the training of heterogeneous models or the handling of diverse tasks [30]. In addition, it is essential for the condensed graph to maintain unbiased representations compared to the original graph, which is particularly crucial in fields sensitive to bias, such as human resources [31], criminal justice [32], and financial services [33]. Moreover, a robust GC procedure can enhance the resilience of condensed graphs to noisy graph structures or features [34] in real-world systems. Beyond these considerations, the scalability of GC also emerges as a key factor in practical applications [35]. An efficient GC method enables rapid adaption in life-long graph learning scenarios, effectively handling the continuous growth and dynamic changes of graph data. These criteria, effectiveness, generalization, efficiency, fairness, and robustness, play a pivotal role in GC evaluation and simultaneously serve as guiding principles for research motivations. Thus, we categorize existing GC methods based on these evaluation criteria, to provide a structured taxonomy for a comprehensive field survey. This taxonomy can help researchers comprehensively understand the diverse motivations driving GC methods, thereby facilitating a clearer path toward achieving their specific objectives.
@@
While recent surveys on related topics including dataset distillation [36, 37, 38, 39], data-centric graph learning [20, 40, 41], as well as graph neural network acceleration [42, 19] have touched upon GC, they summarize GC as a peripheral application or future direction, lacking systematic updates on the latest research and a holistic view. Moreover, concurrent GC surveys [24, 43] and subsequent GC benchmarks [44, 45, 46] focus on detailed optimization strategies, missing high-level perspectives and analysis of design objectives and rationals. This hinders researchers from understanding the novel GC for varying contexts. Recognizing the rapid evolution and growing significance of GC, there is an urgent need for a systematic survey that encapsulates the latest advancements and motivations behind methodological designs.
The contributions of this survey are summarized as follows:
@@
  • We are the first [^1] The initial version was released on January 22, 2024, and is available at https://arxiv.org/abs/2401.11720. to survey the literature on GC and introduce a systematic taxonomy for GC, categorizing existing methods into five distinct groups based on key criteria: effectiveness, generalization, efficiency, fairness and robustness. This taxonomy structures existing knowledge and captures the underlying motivations driving GC research.
@@
  • We provide a comprehensive and up-to-date review of the latest advancements in GC. Additionally, we delve into two fundamental aspects: optimization strategies and condensed graph generation, facilitating an in-depth comprehension of GC techniques.
  • Through rigorous experiments, we systematically compare and analyze diverse optimization strategies in GC based on the five proposed GC evaluation criteria, offering comprehensive guidelines and insights for the design and development of advanced GC methods.
@@
  • We explore practical applications and available resources in GC, while also shedding light on current challenges and future directions in the field. This discussion aims to guide researchers who are seeking to advance GC further.
@@
The structure of this survey is outlined as follows: Section 2 delves into the definition, criteria, and taxonomy of GC. Following this, Section 3 discusses the optimization strategies employed in GC methods. Section 4 provides a thorough investigation of each category of GC, as aligned with our taxonomy. In Section 5, we extensively discuss the methods involved in condensed graph generation, and Section 6 empirically compares the representative GC methods across diverse evaluation criteria. Section 7 is dedicated to summarizing the various applications and resources of GC. Finally, we present the challenges and future directions in this field in Section 8.
@@ 2 Definition and Taxonomy
2.1 Preliminaries
Graph dataset. A large-scale graph dataset can be represented as $\mathcal{T}=\left(\mathcal{V},\mathcal{E}\right)$, which contains $|\mathcal{V}|=N$ nodes and $|\mathcal{E}|=M$ edges  [^2] We unify definitions for node-level and graph-level graph datasets.. Its $d$-dimensional node feature matrix, adjacency matrix, and task specific label are denoted as ${\bf X}\in{\mathbb{R}^{N\times d}}$, ${\bf A}\in\mathbb{R}^{N\times N}$ and $\mathbf{Y}$, respectively. The entry in the adjacency matrix ${\bf A}_{i,j}>0$ denotes an observed edge from node $i$ to $j$, and ${\bf A}_{i,j}=0$ otherwise.
Figure 2: Graph condensation procedure. The original graph and condensed graph are encoded by the relay model $f_{\theta}$ and the condensed graph is optimized according to $\mathcal{L}_{cond}$.
{forest}
for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, line width=0.5pt, draw=black, rounded corners,align=left, minimum width=2em, s sep=5pt, inner xsep=3pt, inner ysep=1pt, , where level=1text width=4.5em, where level=2text width=6em,font=, where level=3font=, where level=4font=, where level=5font=, [Graph Condensation (GC), middle, rotate=90, anchor=north, edge=black [Optimization
Strategy (Sec. 3), middle, edge=black, text width=8em [Gradient Matching, middle, text width=10em, edge=black [GCond [25], GroC [47], HCDC [10], MSGC [48], SGDD [30], FGD [49],
GCARe [50], MCond [26], CTRL [51], DosCond [52], HGCond [53],
EXGC [54], FedGC [55], FedGKD [56], TinyGraph [57], RobGC [58], leaf, text width=29em, edge=black] ] [Trajectory Matching, middle, text width=10em, edge=black [SFGC [28], GEOM [59], GCSR [60], leaf, text width=29em, edge=black] ] [Kernel Ridge Regression, middle, text width=10em, edge=black [KiDD [29], GC-SNTK [61], OpenGC [62], leaf, text width=29em, edge=black] ] [Distribution Matching, middle, text width=10em, edge=black [GCDM [63], CaT [27], GCEM [64], PUMA [65], CGC [66], DisCo [67],
SimGC [68], leaf, text width=29em, edge=black] ] ] [GC Methods
(Sec. 4), middle, edge=black, text width=8em [Effective GC
(Sec. 4.1), middle, edge=black,text width=10em [Optimization-Driven Methods, middle, text width=12em, edge=black [GCond [25], SFGC [28], CTRL [51],
OpenGC [62], GC-SNTK [61],
GEOM [59], KiDD [29], leaf, text width=15em, edge=black] ] [Augmentation-Based Methods, middle, text width=12em, edge=black [MSGC [48], GroC [47], GCSR [60],
PUMA [65], TinyGraph [57], CGC [66], leaf, text width=15em, edge=black] ] ] [Generalized GC
(Sec. 4.2), middle, edge=black, text width=10em [Spectral Methods, middle, text width=12em, edge=black [GCEM [64], SGDD [30], leaf, text width=15em, edge=black] ] [Spatial Methods, middle, text width=12em, edge=black [OpenGC [62], leaf, text width=15em, edge=black] ] ] [Efficient GC
(Sec. 4.3), middle, edge=black, text width=10em [Efficient Graph Encoding, middle, text width=12em, edge=black [GCond [25], PUMA [65], KiDD [29], leaf, text width=15em, edge=black] ] [Efficient Optimization, middle, text width=12em, edge=black [DosCond [52], CaT [27], EXGC [54],
KiDD [29], GCDM [63], PUMA [65],
GC-SNTK [61], Mirage [69], DisCo [67],
SimGC [68], GCEM [64], OpenGC [62],
CGC [66], leaf, text width= 15em, edge=black] ] [Efficient Graph Generation, middle, text width=12em, edge=black [GCond [25], MSGC [48], CGC [66],
GCSR [60], GCDM [63], SFGC [28],
CaT [27], PUMA [65], OpenGC [62],leaf, text width= 15em, edge=black] ] ] [Fair GC
(Sec. 4.4), middle, edge=black, text width=10em [FGD [49], GCARe [50], leaf, text width=29em, edge=black] ] [Robust GC
(Sec. 4.5), middle, edge=black, text width=10em [RobGC [58], leaf, text width=29em, edge=black] ] ] [Condensed Graph
Generation (Sec. 5), middle, edge=black, text width=8em [Feature Generation
(Sec. 5.1), middle, edge=black, text width=10em [Initialization, middle, edge=black, text width=6em [Random, middle, text width=4em, edge=black [GCond [25], DosCond [52], leaf, text width=15em, edge=black] ] [Coreset, middle, text width=4em, edge=black [SFGC [28], GEOM [59], GCSR [60],leaf, text width= 15em, edge=black] ] [Cluster, middle, text width=4em, edge=black [MCond [26], CTRL [51],leaf, text width= 15em, edge=black] ] ] [Closed-Form
Solution, middle, text width=6em, edge=black [CGC [66], leaf, text width=21em, edge=black] ] ] [Structure Generation
(Sec. 5.2), middle, edge=black, text width=10em [Modeling, middle, text width=6em, edge=black [Generative Model, middle, text width=9em, edge=black [SGDD [30], DisCo [67], leaf, text width=10em, edge=black] ] [Parameterization, middle, text width=9em, edge=black [GCond [25], GCSR [60],
KiDD [29], GCEM [64],leaf, text width= 10em, edge=black] ] [Pre-defined Structure, middle, text width=9em, edge=black [MSGC [48], GCond [25],
CGC [66], GCSR [60],
GCDM [63], SFGC [28],
CaT [27], PUMA [65],
OpenGC [62],leaf, text width= 10em, edge=black] ] ] [Sparsification, middle, text width=6em, edge=black [GCond [25], SGDD [30], DosCond [52], KiDD [29], leaf, text width= 21em, edge=black] ] ] ] ]
Figure 3: The taxonomy of graph condensation.
Graph neural network. GNNs leverage the graph topological information ${\bf A}$ and node features ${\bf X}$ to learn node representation via message-passing [13], formulating GNN layers by aggregation and transformation functions. The $l^{th}$ layer of GNNs is formulated as:
$\mathbf{m}_{v}^{(l)}\leftarrow\operatorname{aggregate}\left(\mathbf{h}_{v}^{(l-1)},\left\{\mathbf{h}_{u}^{(l-1)}\mid u\in\mathcal{N}(v)\right\}\right),$ (1)
$\mathbf{h}_{v}^{(l)}\leftarrow\operatorname{transform}\left(\mathbf{m}_{v}^{(l)},\mathbf{h}_{v}^{(l-1)}\right),$
where $\mathbf{h}_{v}^{(l)}$ represents the representation of node $v$ in the $l^{th}$ layer, while $\mathbf{m}_{v}^{(l)}$ denotes the message for node $v$. The function $\operatorname{aggregate}(\cdot,\cdot)$ aggregates the neighboring nodes of node $v$ (i.e., $\mathcal{N}(v)$) to compute the message $\mathbf{m}_{v}^{(l)}$, which is subsequently updated by the transformation function $\operatorname{transform}(\cdot,\cdot)$. Node representations are initialized as node features, and the final representations after $L$ GNN layers are utilized for various downstream tasks.
@@ 2.1 Preliminaries
Graph condensation. Graph condensation aims to find a small condensed graph dataset $\mathcal{S}=\left(\mathcal{V}^{\prime},\mathcal{E}^{\prime}\right)$ with $|\mathcal{V}^{\prime}|=N^{\prime}$ and $N^{\prime}\ll N$, enabling more efficient GNN training while maintaining performance comparable to GNNs trained on the original graph $\mathcal{T}$. The adjacency matrix, node feature matrix, and task-specific label of $\mathcal{S}$ are denoted as ${\mathbf{A}^{\prime}}$, $\mathbf{X}^{\prime}$ and $\mathbf{Y}^{\prime}$, respectively. The compression rate (a.k.a. condensation ratio) $r$ is defined as $\frac{N^{\prime}}{N}$. To facilitate the connection between original graph $\mathcal{T}$ and condensed graph $\mathcal{S}$, a relay graph model $f_{\theta}\left(\cdot\right)$, parameterized by ${\theta}$, is employed in the optimization process for encoding both graphs as shown in Figure 2. Then, the graph condensation is formulated as an optimization problem:
$\mathcal{S}=\arg\min_{\mathcal{S}}\mathcal{L}_{cond}\left(f_{\theta}\left({\mathcal{T}}\right),f_{\theta}\left({\mathcal{S}}\right)\right),$ (2)
@@ 2.2 Criteria for Graph Condensation
The primary metric for assessing GC methods involves evaluating the accuracy of GNN models trained on condensed graphs as compared to those trained on the original graphs, providing a direct measure of information preservation. To achieve a comprehensive evaluation and gain a holistic understanding of method performance, it is essential to incorporate additional criteria. Specifically, we summarize the key criteria of GC as follows:
  • Effectiveness assesses the accuracy of a specified GNN architecture trained on the condensed graph under various compression rates [25]. It measures the extent to which the condensed graph preserves the essential characteristics of the original graph.
  • Generalization measures the overall accuracy of different GNN architectures trained on the condensed graph across various downstream tasks [30]. It reflects the condensed graph’s adaptability and utility in diverse contexts.
  • Efficiency assesses the time required to condense a graph [52], focusing on the speed of condensed graph generation and the method’s feasibility for time-sensitive scenarios.
  • Fairness evaluates the disparity in model performance between GNNs trained on condensed graphs and those trained on original graphs across different demographic groups[49], aiming to prevent bias amplification in the condensed graph and ensure equitable model performance.
  • Robustness measures the quality of condensed graphs generated from the noisy original graph. It assesses the ability of GC methods to identify and preserve the vital characteristics of the original graph despite the presence of noise in structures or features.
Given these criteria, an ideal GC method should generate a high-fidelity condensed graph dataset at low cost, enabling models with different GNN architectures to achieve results comparable and unbiased to those obtained on the original graph on a wide range of downstream tasks.